Oxford Materials IT Committee Meeting

A meeting of the Department of Materials IT Committee was held on Tuesday 24th October 2006.


  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters arising from the Minutes of the Meeting of the last meeting
  • 3 IT/AV – Matters for Discussion (attached)
  • 4 IT/AV – Head of IT Report (attached)
  • 5 University IT Users Group – brief verbal report (PJW)
  • 6 IT/AV expenditure - matters for report (PJW)
  • 7 AOB
  • 8 Date and time of next meeting: Tuesday 30th January 2007 at 2pm.


See minutes below.


Oxford Wireless LAN - KS to update on progress over this summer and current status

Websites - PJW to update on progress over this summer and current status

Server provision - AMC to update on progress over this summer and current status.


Admin Network
New Begbroke finance officer Laura Vockins was added to the system. APD's printer was replaced. Extra disks were added to the server due to increased storage requirements.

Teaching Laboratory
No changes to report. Several laptops were loaned out over the summer but returned ready for the start of term.

Lecture theatres / meeting rooms
The 21 BR lecture theatre was refurbished in September. Some problems with the initial specification have been rectified by Adrian Taylor and the new equipment is now fully functional. AHMcK and PJW have been liasing with Ian Campbell and Robert Taylor in Physics to learn about the new Niagra Viewcast system purchased for divisional use. The weekly EM seminars and Departmental Seminars are both being broadcast live each week this term, primarily for remote audiences at Begbroke although other users around the department have been accessing the system too. Further work is needed to learn how to do post-processing of the webcast and where to host the data for anytime access on the web. Overall webcasting is very time intensive with a high cost to the department.

Digital Darkroom
Intermittent smearing problems with the HP8550 colour printer continue to be fixed by Andrew McKnight (once per month). The second PC was swapped with the almost identical 21BR lecture theatre PC.

Nothing to report.

The MML was completed relocated for one month during re-wiring works in 21BR. The new UPS system was installed and the network upgraded to gigabit.
The 48q cluster continues to have hardware problems on two nodes (8q) but otherwise has been fully utilized. Plans for expansion of the cluster are being costed and are likely to occur before Christmas. However the cluster will need to be relocated to avoid over heating problems.

Nothing to report.

The department website has been updated for the new academic year (mostly). Research in Progress (RIP) is still being collated prior to staff homepages being updated.

PJW spent several weeks over the summer getting to grips with PHP and MySQL by porting the RIP database and the EM booking database to the departmental linux webservers (http://external.materials.ox.ac.uk and http://internal.materials.ox.ac.uk). The changeover is scheduled to occur at Christmas and the old Mac server (http://photos.materials.ox.ac.uk) running Filemaker Pro 2.1 will be retired.

PJW has also been expanding the RIP database system into a web-based database system for tracking users, rooms, phones, and other core departmental data, aiming to automate as much as possible and reduce re-typing and duplication of data.

PJW attended the OUCS workshop on Content Management Systems(CMS). OUCS are pursuing OpenCMS as a technical solution and PJW has started to investigate it, but is not yet in a position to say whether it is suitable solution for the department.

Meanwhile OMCS have commissioned a BegbrokeNano website which will use the Joomla CMS.

The intended review of research group websites last summer still has not happened because of time-limitations and taking a broad departmental view rather than fixing specific research group websites individually.

The AIK group website http://www-eia.materials.ox.ac.uk was developed over the summer by Grigore Moldovan and can act as a good example of research group website. The DJHC group intends to use the similar system.

A department handbooks were updated for the new academic year and printed copies circulated. (separate summary to be circulated).

The hot summer weather and more particularly the high heat output of the new cluster caused the temperature in the HR plant room to become excessively high. This is thought to have been the cause of the intermittent failures of the firewall over the summer. An alternative location for housing high-heat-output cluster(s) is needed within the next six months, especially since the MML plans to expand the cluster capabilities.

OUCS completed installation of new FroDo network cabinets into the buildings HR, ETB, HB, BR and BB. PR is just waiting for delivery of a smaller network cabinet due to space contraints. The FroDo equipment provides network services to each building. The Frodo cabinets are being used to provide Oxford Wireless LAN within the department. (see separate report).

After considerable discussion with OUCS the department will continue to use the existing fibre cable infrastructure between buildings for our core network provision independent of FroDO. This has enabled us to commence the upgrade of the links between buildings to gigabit for £700/building, without incurring the £5k/building OUCS backbone charge. The link HR-BR is now running at Gigabit. The link HR-HB will be upgraded next, possibly followed by the link HR-BB. The other links HR-PR and HR-ETB do not carry sufficient traffic to bother upgrading at present. Upgrading the backbone connection to Gigabit is deemed not necessary at present, although the next firewall currently under development will have gigabit capability. The MML now has Gigabit networking, but the majority of the department has 100Mbit. The department still has eight 10Mbit switches in operation now which should be phased out over the coming year by replacing core switches and trickling older switches.

Numerous network switches were reconfigured and moved around as part of the general network upgrade works to implement Oxford Wireless LAN within the department (see separate report).

Rewiring of BR was completed in August.

Materials2(main netware fileserver) storage capacity was increased by 72Gb
Oums-nomad (Altiris Deployment Server, Sophos EM Library, FM Group database, NUTD) now also serves Now Up-to-date.
Oums-wsus (windowsupdates, Sophos EM console) server is working well.
Oums-sus(the old windowsupdates) server now only has win98 antivirus connections.
www-int, www-ext (webservers) continue to function well.
Databases(the old sophos) was hacked and taken off line and is being investigated by OUCS.
Installations using Novell OES Linux and Novell OES Netware are being run on Materials1 and ittemp3.
Materials0 was reinstalled, the disk capacity expanded to 3Tb and is now ready for use. Users are being transferred from emgroup1 prior to decommissioning emgroup1.

APL was supplied with the phone from the Yarnton Meeting room. The write-up room (and Rashid Bhatti) were without a phone for several months due to the power transformer being misplaced. There have been very few reports of problems with the phones.

One computer was hacked (server databases.materials which was about to be decommissioned anyway.). Several instances of possible p2p activity were investigated but not action taken. The door entry system in PR has an intermittent hardware fault. We have a temporary workaround but it needs fixing in the longer term.


A meeting of the Department of Materials IT Committee was held on Tuesday 24th October 2006.
The following people attended: APD, PJW, AMC, KS, JAMS, CJS, and Ralf Drautz.

1. Apologies for Absence and Membership of the Committee

AHM off sick.

JAMS term of office expired in TT2006, however JAMS agreed to be reappointed until TT2008.

2. Matters arising from the minutes

a) Email and shared calendar server (Minute 2a)
Requirements of an Exchange server have been investigated. Informal comments on Lotus Notes as used in Central Admin were unfavourable. Other systems were considered but all had pros and cons with there being no clear optimum system. Groupwise remains of interest but current needs for shared calendaring do not justify the time necessary to implement it.

Users requiring calendar sharing are recommended to use Now-Up-To-Date software which is managed by Department IT staff. This software is a free-standing solution not integrated with email client software. Cost of Now-Up-To-Date is around $100/machine perpetual.

Users requiring calendar sharing and integration with Outlook are advised to use the new pay-per-use Exchange server system run by NSMS(OUCS). Cost is around £100/user/year.

Groups of users of Now-Up-To-Date are: Admin=APD+AJJ+CS, DJHC+KH, OMCS=JAAC+CJS+FC+KJ+MR+JPN (trial uptake), Faraday=CJ+RY+LJ (possible interest)

b) IT disaster recovery plans (Minute 2b)
PJW circulated, with the agenda, a draft review of UPS provision within the department as the first step towards creating disaster recovery documentation. The department has spent several thousands of pounds over the past year improving UPS cover for servers and network switches but further work is needed in this area so the item is still carried to the next meeting. [Carried to next meeting.]

Action: PJW

c) Website (Minute 2c)
The proposed review of research group websites did not happen during the summer but has been carried forward to the next meeting. Advice and guidance is to be offered by PJW upon compliance with University Website Policies and upon mechanisms for easy updating of content. PJW to email notices to raise interest in web authoring and best practice. Carried to next meeting. [Item carried to next meeting however there has been considerable groundwork over the summer as outlined in the Head of IT report]

Action: PJW

d) Copyright (Minute 2d)
PJW attended lecture 'WISER Teaching: Copyright Law' in June to gain overview of the new digital copyright license regulations. On 15th August the department registered a 6-monthly 'null' return. In October PJW notified seniorstaff, fellows and admin by email about the implications of the HE digital copyright license to cover the creation of digital copies for use a course material available electronically.

This department does not have a centrally coordinated provision of course material electronically. Most lecturers give paper handouts. Some lecturers make their course materials available electronically themselves.

PJW to email staff to promote use of weblearn to coordinate distribution of course materials.

Action: PJW

e) Website (Minute 2e)
JAMS raised several points concerning the department website. The new projects page needs to be clearer for prospective students to find projects of interest. The current status of representation of research groups on the Department website is patchy at best, and a centralized system for summarizing Research Group information is desirable. PJW held a first meeting with JAMS and is currently drafting a template for a research group website to be rolled out over the summer. [Item carried to next meeting however there has been considerable groundwork over the summer as outlined in the Head of IT report]

Action: PJW/JAMS

f) Computer misuse procedures (Minute 4)
The new procedures for dealing with computer misuse by students (formulated by AOT in response to a repeat offence of p2p filesharing) were presented to DC on Monday 12th June and formally adopted. These procedures were added to the Facilities and Procedures handbook in Appendix B following on from Appendix A which covers the department's IT Rules and Regulations. These procedures were also added to the Graduate Handbook in Section 23. The Undergraduate Handbooks do not contain these procedures since undergraduates are subject to discipline via their colleges but do contain a full copy of the University Rules in Appendix E.

g) Microwave (Minute 7a)
APD asked for an update on the status of the microwave equipment. The equipment has been turned off and the license discontinued but there has been no progress in finding a buyer. PJW should sell or dispose of the equipment. KS reported that a buyer for the microwave could not be found so the equipment will be removed and disposed.

Action: PJW/KS

3 Matters for Discussion

KS gave a verbal report on the many network changes that have occurred over the summer and in particular the current status of implementation of the Oxford Wireless LAN [five access points being installed in communal areas, three more locations identified; cost £300/access point]. The written report was circulated with these minutes.

There was discussion about the effect upon the department of the various network outages over the summer due to several hardware failures partly brought about by the necessary configuration changes to support OWL. The strategy for hardware upgrades to the network was discussed and the rolling upgrade programme for switches was outlined.

It was noted that every effort was made to schedule planned disruptions to low usage times (e.g. weekends in the summer). Unfortunately some changes although successful at the time introduced longer term instabilities which came to light several days later.

There was discussion about how to improve the communications between users and IT staff during outages. The existing system of users reporting by phone to Reception and IT staff reporting by tannoy remain the best practical methods for Oxford site. However at Begbroke it was proposed that IT staff should use Begbroke reception staff to update the whiteboard at the entrance to Hirsch Building since there is not a tannoy system at Begbroke.

IT staff have already implemented increased monitoring of network switches to provide faster and automated notification of problems, however IT staff still require that members of department must to report problems to Reception 73777 giving details.

a) CJS reported that some VOIP phones reboot automatically intermittently. PJW will investigate whether remote monitoring can detect such problems.

Action: PJW/KS

PJW outlined the progress over this summer and current status of the various website projects as described in the Head of IT report circulated with the agenda.

Discussion of server provision was postponed to next meeting.

4. Head of IT Report

A written report was circulated with the agenda (above). Key points were highlighted.

a) PJW to email notices with details about the new equipment available in the 21BR lecture theatre.

Action: PJW

b) This term the EM and Departmental seminars are being webcast live, primarily for Begbroke users to view the seminars. Archiving of seminars is also being investigated. PJW to present summary document to DC in week 8 overviewing equipment/personel costs for webcasting. The committee also discussed possible legal issues, possible uses for academics etc.

Action: PJW

c) Printing problems in the digital darkroom were discussed and it was agreed to fund a replacement for the hp8550a3 colour printer.

Action: PJW

5. Oxford University IT Users Group Meeting http://www.ict.ox.ac.uk/oxford/groups/ouitug/

a) ITUG is being replaced by the ICT Forum as part of the new ICT strategy.

b) IT staff to circulate to ITC a summary of issues from the new University ICT Strategy.

Action: PJW/AMC/KS

c) IT staff to circulate to ITC a summary of issues from the new OUCS 5-year plan.

Action: PJW/AMC/KS

6. IT/AV Expenditure

A breakdown of the expenditure for last year was distributed. The total expenditure for the financial year 1st August 2005 - 31st July 2006 was £41k. A categorised breakdown of expenditure was also circulated and discussed in terms of financial planning.


a) The function of the committee includes advising the Department upon IT strategy. PJW requested all members of the committee to take some time to consider the wider picture and provide some feedback by email. Possible areas of interest raised instantly included: department-wide data archive procedures, email storage/archiving given limited disk space on herald, automating use of idle computational resources, increased use of skype.

Action: ALL

b) PJW to update committee webpage http://www.materials.ox.ac.uk/it/itcommitee.html to take account of revised committee structures within the department.

Action: PJW

8 Date of Next Meeting

The next meeting will take place on Tuesday 30th January 2007 at 2pm in 21BR Committee Room (room booked already)

The meeting closed at 3.55pm